OpenStudy (anonymous):

For the right triangle shown, the lengths of two sides are given. Find the third side. Leave your answer in simplified, radical form. a = 5, b = 10, c =

OpenStudy (linn99123):

c = sqrt(a^2 + b^2) =13

OpenStudy (anonymous):

13!?

OpenStudy (anonymous):

\[c^2=5^2+10^2=25+100=125\]